OpenStudy (anonymous):

find the center and the radius of the circle with this equation x^2+y^2+4x+2y+3=0

OpenStudy (anonymous):

\[x ^{2}+y ^{2}+4x+2y+3=0\]

OpenStudy (anonymous):

the center is (-2,-1) yes?

OpenStudy (anonymous):

U need to get this into the form (x-a)^2 + (y-b)^2 = r^2

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so the radius is either sqrt2 or sqrt3\[\sqrt{2} or \sqrt{3}\]

OpenStudy (anonymous):

Now just shuffle things around till u get an integer on the right and the radius is the positive square root.

OpenStudy (anonymous):

is it the sqrt3?

OpenStudy (anonymous):

I haven't done it, let me look..

OpenStudy (anonymous):

I make it sqrt 2

OpenStudy (anonymous):

Your (x+2)^2 gives you a -4 and your (y+1)^2 gives you a -1 so -4-1 +3 = -2 is 2 on the other side, take square root.